Consider a bond paying a coupon rate of 12.25% per year semiannually when the market interest rate is only 4.9% per half-year. The bond has six years until maturity.

a. Find the bond's price today and six months from now after the next coupon is paid. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Current Price: 1109.19 (This is the answer)

Price After 6 Months: (Answer this Please)

b. What is the total rate of return on the bond? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Please calculate % per 6 months

Total Rate of Return: (Answer this)
Answer #1

Price after 6 months

Using financial calculator
Input:

Fv = 1000

N = 6*12 -1 = 11

PMT = 12.25%*1000/2 = 61.25

I/Y = 4.9

Solve for PV as -1102.29

Price after 6 months = $1102.29

b: Total rate of return = (Closing price-Opening price+Coupon)/ Initial price

= (1102.29-1109.19+61.25)/1109.19

= 54.35/1109.19

= 4.9%